opPlaskett's binary system consists of two stars tha
Answer #1
The idea is to use two concepts:
circular motion:     F = m v2 /r
and gravitational force:    F = G mm / d2    
Notice that in this case,   d =2r     so we can write
      G m m / (2r)2 = m v2 / r
        G m /4r = v2         
but what is r?      well, you aregiven the speed and time time to orbit once, so...
   time = 12.5 days = 1080000 seconds
   speed =   240000 m/s
distance for one orbit is    2pi r      which isalso    speed * time    so
     2pi r = 240000 *1080000
       r = 4.127 x1010 meters
And now...
   m = 4 r v2 / G  = 4 * 4.127 x 1010 * 2400002 / 6.67 x10-11 =
        =   1.425x 1032   kg
Our sun has mass    1.99 x1030 kg so this star is
    142.5x 1030 / 1.99 x1030 =    71.64 timesmore massive than our sun
Answer #2

Orbital period = 12.5*24*60*60 s = 1080000 s

Speed = 240 km/s


So, R = V*t / 2*pi = 41252961249.42 m


So, M*v^2/R = GM^2/(2R)^2


So, M = 4(v^2)R/G = 1.425*10^32 kg
Answer #3

G*M^2/(2R)^2=Mv^2/R


==> M=4v^2*R/G


and R=v*t/2*pi